# Abundant Population of Broad H Emitters in the GOODS-N Field Revealed by CONGRESS, FRESCO, and JADES

### v1

@@ 5.3 Are Broad Line Emitters LRDs?
In this section, we evaluate the fraction of our spectroscopically confirmed broad H emitters that can be photometrically classified as LRDs by applying the selection criteria outlined in Rinaldi etĀ al. (2024):
ā¢
@@
The left panel of Figure 8 shows the positions of the broad H emitters on the color-color plot. We find that a significant fraction of our sample (42%) cannot be photometrically selected as LRDs, with the majority (7 out of 8) failing to meet the red slope criterion. This result is consistent with Hainline etĀ al. (2025), who applied a stricter red slope threshold and found that only 30% of spectroscopically identified BLAGN with JWST meet the LRD selection criteria. These findings suggest that not all broad-line emitters with a compact, red appearance exhibit the characteristic āV-shapedā SED of LRDs, and that many lack a steep red slope.
Additionally, since bright H emission is observed in the F444W for all FRESCO sources, applying a relatively loose red slope criterion raises concerns about the FRESCO sources that satisfy the LRD color selection. It remains unclear whether these sources intrinsically exhibit such steep red slopes, or whether their observed colors are primarily driven by line-boosting effects (Hainline etĀ al., 2025).
A more conservative approach to selecting LRDs with a high AGN yield through color cuts involves applying a higher cut on the red slope, in order to reduce contamination from line-boosting effects, as proposed by Barro etĀ al. (2024) and Greene etĀ al. (2024). In the left panel of FigureĀ 8, we add the selection criteria proposed in both studies:
ā¢
Red slope: F277W F444W 1.5 mag (Barro etĀ al., 2024)
ā¢
Red slope: F277W F444W 1.6 mag (Greene etĀ al., 2024)
@@
In addition to color-based selection, Kocevski etĀ al. (2024) proposed an alternative approach for identifying LRDs by applying cuts on the UV and optical continuum slopes, derived from fitting photometry across multiple bands blueward and redward of the Balmer break at 3645Ć . This method enables the detection of LRDs with reduced contamination from galaxies that exhibit strong Balmer breaks but lack a rising red continuum. Given the lack of full photometric coverage for some sources, we derive their UV and optical continuum slopes using F090W+F115W+F150W and F277W+F356W+F444W, respectively, to make optimal use of the available photometry. For sources lacking photometry in F150W or F277W, we estimate their fluxes in these bands from their best-fit SEDs.
We adopt the same cuts used in Kocevski etĀ al. (2024):
ā¢
@@
The right panel of FigureĀ 8 shows the distribution of our sample in the plane. Based on this selection approach, 74% of our sample do not satisfy the selection criteria. Similarly, all of the outliers fail to satisfy the red-slope criterion. All three LRD selection techniques miss an important fraction of our broad-line emitters (i.e., 42%, 84%, and 74%), with two of them failing to identify the majority, suggesting that current LRD selection methods may overlook a substantial fraction of the broad-line emitter population.
@@
However, most of the non-LRD broad H emitters do not satisfy the red slope criterion, with the majority originating from CONGRESS. In these cases, the [OĀ III] doublet falls within the F277W filter, making the F277W F444W color less red, and causing these sources to fall outside the selection boundary. An example of an object that clearly exhibits LRD-like behavior in its SED but is not color-selected as an LRD due to strong [OĀ III] emission is presented in Rinaldi etĀ al. (2025). This further highlights the challenges and limitations of color-based LRD selection at these redshifts.
@@
While several alternative color selection criteria of LRDs have been proposed in the literature (e.g., Greene et al. 2024; Pérez-GonzÔlez et al. 2024), we adopt the criteria used by Rinaldi et al. (2024) in this work to maintain consistency, as their LRD sample is also constructed in the GOODS-N field.
@@
To further explore the relationship between broad-line emitters and photometrically selected LRDs, we perform a comparison between the average SEDs of our broad H emitters and the LRD sample selected in Rinaldi etĀ al. (2024). Here, we only include LRDs with available spectroscopic redshifts. The average SED of the LRD sample is derived using all available photometric data and following the methodology outlined in PĆ©rez-GonzĆ”lez etĀ al. (2024). In contrast, the average SED of our broad H emitter sample is constructed using the best-fit SEDs obtained in Section 5.2, as about half of the sources in our sample have photometry available in only 6 filters. Figure 9 displays the average rest-frame SEDs for both broad H emitters and LRDs, with flux normalized at 0.4ām. The average SED of LRDs selected in the GOODS-S field presented in PĆ©rez-GonzĆ”lez etĀ al. (2024) is also included for comparison.
@@
Overall, we do not observe significant differences between the average SEDs of our broad H emitters and LRDs within the rest-frame 0.1ā1ām range. The SEDs closely align up to rest-frame 0.4ām, suggesting that LRDs may share similar UV emission properties with broad-line emitters. However, beyond 0.5ām, the average SEDs of LRDs exhibit a noticeably steeper slope toward longer wavelengths compared to those of broad-line emitters, with particularly significant deviations at 0.5 and 0.65ām. Given that the observed continuum in the average LRD SEDs between these wavelengths is comparable to that of broad-line emitters, we suspect that the steep red slope is not an intrinsic feature of at least some LRDs, but rather a result of line-boosting effects from strong emission lines, such as [OĀ III] and H, falling within individual wide filters (Hainline etĀ al., 2025). Figure 9 further supports the concern that adopting a low red slope cut for LRD selection may lead to contamination by sources that exhibit āfakeā steep red slopes induced by such line-boosting effects, particularly when there are strong emission lines falling within the filters used for color selection.
5.4 Detectability of Broad H Line in Grism Spectra
@@
We identify 19 photometrically selected LRDs from Rinaldi et al. (2024) that lie within the FRESCO and CONGRESS fields and have H emission detected in the grism spectra. However, 68% of these sources (13 out of 19) do not show broad H line profiles (i.e., FHWM 1000 km/s), with the exception of the 6 broad H emitters identified in this work. Such a low fraction of LRDs displaying broad-line features is consistent with the findings reported by Pérez-GonzÔlez et al. (2024).
@@
Although no broad-line feature is observed in these LRDs, we cannot fully rule out the possibility that they are broad-line emitters, as the broad component could be hidden due to the insufficient sensitivity of the grism spectra. Notably, two sources that show no broad line feature in the grism spectra have been reported to exhibit broad line features in NIRSpec medium-resolution grating data in previous work (Maiolino etĀ al., 2024). To illustrate how S/N may affect the detection of broad components, we present both the NIRSpec/grating and NIRCam/grism spectra for these two sources in Appendix C.
@@
Figure 10 presents the distribution of our broad-line emitters and the photometrically selected LRDs of Rinaldi etĀ al. (2024), which lack broad-line features in the grism spectra, on the redshift-magnitude plane. Overall, our broad-line emitters are significantly brighter than the LRDs, with a division at 26 mag that effectively separates the two populations. Notably, the two sources that exhibit broad-line features in the grating spectra but not in the grism are both fainter than 26 mag, suggesting that some faint LRDs may indeed be broad-line emitters, but their broad components are too faint to be detected in grism observations. In other words, the broad-line search with the CONGRESS/FRESCO data is effective only down to ā26āmag.
@@ 6 Conclusion
ā¢
Among the 19 broad H emitters, 18 of them have a broad component that contributes more than 50% of the total H flux. Additionally, [OĀ III]Ā emission line is detected in three sources, with line widths significantly narrower than those of the broad H components, suggesting that the observed broad features of these sources are mainly driven by the AGN BLR.
ā¢
@@
ā¢
The stellar masses of the host galaxies in our sample were derived through SED-fitting using Prospector. Consistent with previous studies, our sample exhibits higher ratios relative to the local relations reported by Reines & Volonteri (2015) and Greene etĀ al. (2020). However, given that our sample is selected based on the detection of broad H emission, some degree of selection bias may be present. In particular, low sources may be missed due to the limited sensitivity of the grism, which could prevent the detection of faint broad-line components.
ā¢
By testing various LRD selection techniques proposed in the literature, we find that all of these methods will miss a significant fraction of our spectroscopically confirmed broad H emitters, while the majority of the non-LRD broad H emitters lack the steep red slopes. This suggests that not all broad-line emitters with LRD-like properties exhibit the characteristic āV-shapedā SEDs of LRDs.
ā¢
To further investigate the nature of LRDs, we compared the average SED of our broad-line emitter sample with that of the photometrically selected LRDs from Rinaldi etĀ al. (2024) in the same field. The average SEDs are largely consistent with each other, except for a steeper red slope in the LRD sample at rest-frame wavelength 0.5ām, likely due to the contribution of [OĀ III] and H emission. Given the much lower continuum levels observed between these two lines, we conclude that the red slope, at least in some LRDs, would be naturally less steep if emission lines are removed, as discussed in Hainline etĀ al. (2025).
ā¢
We further examined the LRDs in the GOODS-N field with H detections in the grism spectra and find that 68% do not exhibit broad H features. This suggests that the LRD population may not only consist of broad-line emitters. Additionally, we identified two LRDs that lack broad H features in the NIRCam grism spectra but do show broad components in the NIRSpec grating spectra. Both of these sources are fainter than 26 mag in the filter where H is observed, whereas the majority of our broad H emitters are brighter than this threshold. This implies that the broad-line search with the CONGRESS/FRESCO data is effective only down to ā26āmag, and therefore that some LRDs may indeed host broad-line components that are undetected in the grism data due to its limited sensitivity.
@@ Acknowledgments
EE, YZ acknowledge support from the NIRCam Science Team contract to the University of Arizona, NAS5-02015. F.W. acknowledges support from NSF award AST-2513040. AJB acknowledges funding from the āFirstGalaxiesā Advanced Grant from the European Research Council (ERC) under the European Unionās Horizon 2020 research and innovation program (Grant agreement No. 789056). RM acknowledges support by the Science and Technology Facilities Council (STFC), by the ERC through Advanced Grant 695671 āQUENCHā, and by the UKRI Frontier Research grant RISEandFALL. RM also acknowledges funding from a research professorship from the Royal Society. BER acknowledges support from the NIRCam Science Team contract to the University of Arizona, NAS5-02015, and JWST Program 3215. ST acknowledges support by the Royal Society Research Grant G125142. GV acknowledges support by European Unionās HE ERC Starting Grant No. 101040227 - WINGS. The research of CCW is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
@@
This work is based on observations made with the NASA/ESA Hubble Space Telescope and NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1181 (JADES), 1895 (FRESCO), and 3577 (CONGRESS). The specific JWST observations analyzed can be accessed via https://dx.doi.org/10.17909/8tdj-8n28 (catalog doi:10.17909/8tdj-8n28) and http://dx.doi.org/10.17909/6rfk-6s81 (catalog doi:10.17909/6rfk-6s81 ). Support for program #3577 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. The authors acknowledge the FRESCO team for developing their observing program with a zero-exclusive-access period.
Appendix A Non-Confirmation of Two Broad H Emitters Reported by Covelo-Paz et al.Ā (2024)
@@
In our sample, we exclude two broad H emitters reported by Covelo-Paz etĀ al. (2024) due to concerns regarding the validity of their classification. For the first source (JADES ID: 1080661; RA = 189.27616 deg, Dec = 62.21416 deg), the flux ratio does not meet our compactness criterion. FigureĀ 11 presents the 1D and 2D grism spectra obtained in F356W from CONGRESS. The F356W direct image reveals an extended morphology for this source. We also performed a 1D line fit of its H line profile. As illustrated in FigureĀ 12, even after accounting for potential morphological broadening, the best-fit FWHM is only 500 km/s, which does not satisfy our requirement for the line width of a broad H emitter. Additionally, we note that our best-fit FWHM differs from the value reported in Covelo-Paz etĀ al. (2024) (i.e., 1674 km/s). However, the cause of this discrepancy is unclear.
@@
For the second source (JADES ID: 1089616; RA = 189.18713 deg, Dec = 62.27289 deg), the top panel of FigureĀ 13 shows its 1D and 2D grism spectra in F356W from CONGRESS. While a strong emission feature is detected at 3.8 m, we observe a noticeable spatial offset between the position of the emission line and that of the source in the direct image, raising concerns about its association with the target. In the bottom panel of FigureĀ 13, which presents the F444W grism spectra from FRESCO for the same source, an emission line at 4.15 m is detected with no spatial offset and a S/N4. Combined with the detection of another emission line at 3.17 m, we identify these features as H and [OĀ III] of this source, respectively, confirming the redshift of the source is 5.324. The relatively weak line strengths are also consistent with the faint nature of the source in the direct images. Upon further inspection, the strong 3.8 m line is confirmed to be [SĀ III] from a nearby bright source (JADES ID: 1029833; RA = 189.20103 deg, Dec = 62.26567 deg). Figure 14 presents the 1D and 2D grism spectra of this source in the F356W and F444W filters. The [SĀ III] emission line of this source matches the 3.8Ā m emission line observed in the 2D spectrum of source 1089616 very well, suggesting the 3.8Ā m emission line is not H of 1089616, but [SĀ III] of 1029833.
@@
Appendix B Discrepant Broad-Line Widths for a BLAGN Reported by Greene et al.Ā (2024)
@@
Recently, the JWST Cycle 3 General Observer program Slitless Areal Pure-Parallel High-Redshift Emission Survey (SAPPHIRES-PID: 6434; PI: E. Egami; Sun etĀ al. 2025a) revisited the Abell 2744 field and obtained NIRCam grism spectrum (R1600) for one of the BLAGN (Prism ID: 38108; SAPPHIRESās ID: 39082) identified in Greene etĀ al. (2024). Figure 15 shows the observed H line in 1D and 2D grism spectra obtained in F444W filter, along with the line fitting results. The 1D grism spectrum reveals a broad H component with a line width of 1300 km/s, which contrasts significantly with the prism-based value of 4000 km/s. For comparison, we also include the H broad component observed in prism by using the FWHM and H flux (treated as an upper limit of the flux for the broad component) reported by Greene etĀ al. (2024). Even a modest difference in the FWHM measurements can result in a substantial variation in the inferred . In Figure 5, we plot the AGN bolometric luminosity and black hole mass derived for this source from the SAPPHIRES NIRCam grism data (the āXā marker). With the smaller inferred from the grism spectrum, the source becomes more consistent with the distribution of our sample, providing a potential explanation for the discrepancy in black hole mass estimates.
@@
However, the underlying cause of the discrepancy in FWHM measurements remains unclear. Taking into account the large uncertainty associated with the prism-based FWHM, our measurement and that from Greene etĀ al. (2024) are broadly consistent with each other at the 2- level. Since our line fitting was performed on spectra without continuum subtraction, we can rule out the possibility that the broad component was lost due to over-subtraction of continuum. There are two possible explanations for this discrepancy. First, the difference in FWHM values may arise from differences in the respective LSFs. Second, the grism observation may lack the sensitivity required to detect such a faint broad component. Furthermore, we note that Greene etĀ al. (2024) identified broad-line sources using slightly different selection criteria, in particular requiring the FWHM of the broad H component to 2000 kmĀ s-1. This difference in criteria may also contribute to the discrepancy in FWHM measurements.
@@
In this work, we identify two photometrically selected LRDs from Rinaldi etĀ al. (2024) that exhibit broad H line profiles in the NIRSpec medium resolution grating spectra presented by Maiolino etĀ al. (2024), but not in the NIRCam grism spectra, likely due to the limited sensitivity of the grism observations. Figure 16 illustrates the differences in the observed H line profiles for these two sources between the grism and grating data. The S/N of the grating spectra are slightly lower than those presented in Maiolino etĀ al. (2024), possibly due to differences in the spectra extraction methods. In both cases, the H lines appear significantly narrower in the grism spectra compared to the grating spectra. Additionally, the grism continuum exhibits a much higher noise level, supporting the possibility that the broad components are undetected due to the lower S/N of the grism observations.

### v2

@@ V.4 Are Broad Line Emitters LRDs?
In this section, we evaluate the fraction of our spectroscopically confirmed broad H emitters that can be photometrically classified as LRDs by applying the selection criteria outlined in Rinaldi et al. (2024):
@@
The left panel of Figure 9 shows the positions of the broad H emitters on the color-color plot. We find that a significant fraction of our sample (42%) cannot be photometrically selected as LRDs, with the majority (7 out of 8) failing to meet the red slope criterion. This result is consistent with Hainline et al. (2025), who applied a stricter red slope threshold and found that only 30% of spectroscopically identified BLAGN with JWST meet the LRD selection criteria. These findings suggest that not all broad-line emitters with a compact, red appearance exhibit the characteristic “V-shaped” SED of LRDs, and that many lack a steep red slope.
Additionally, since bright H emission is observed in the F444W for all broad H emitters identified in FRESCO, applying a relatively loose red slope criterion raises concerns about those sources that satisfy the LRD color selection. It remains unclear whether these sources intrinsically exhibit such steep red slopes, or whether their observed colors are primarily driven by line-boosting effects (Hainline et al., 2025).
A more conservative approach to selecting LRDs with a high AGN yield through color cuts involves applying a higher cut on the red slope, in order to reduce contamination from line-boosting effects, as proposed by Barro et al. (2024) and Greene et al. (2024). In the left panel of Figure 9, we add the selection criteria proposed in both studies:
Red slope: F277W F444W 1.5 mag (Barro et al., 2024)
Red slope: F277W F444W 1.6 mag (Greene et al., 2024)
@@
In addition to color-based selection, Kocevski et al. (2024) proposed an alternative approach for identifying LRDs by applying cuts on the UV and optical continuum slopes, derived from fitting photometry across multiple bands blueward and redward of the Balmer break at 3645Å. This method enables the detection of LRDs with reduced contamination from galaxies that exhibit strong Balmer breaks but lack a rising red continuum. To make optimal use of the available photometry, we derive the UV continuum slope using the F090W, F115W, and F182M bands. For sources at redshift , we use F210M, F277W, and F356W to measure the optical continuum slope, while for higher-redshift sources (i.e., ), we instead use F277W, F356W, and F444W. For sources lacking photometry in F277W, we estimate their fluxes in F277W from their best-fit SEDs. The uncertainties on the measured slopes are estimated using a Monte Carlo approach following the methodology described in Kocevski et al. (2024).
We adopt the same cuts used in Kocevski et al. (2024):
@@
The right panel of Figure 9 shows the distribution of our sample in the plane. Based on this selection approach, 47% of our sample do not satisfy the selection criteria. Similarly, most of the outliers fail to satisfy the red-slope criterion. All three LRD selection techniques miss an important fraction of our broad-line emitters (i.e., 42%, 84%, and 47%) suggesting that current LRD selection methods may overlook a substantial fraction of the broad-line emitter population.
@@
However, most of the broad H emitters that failed the LRD color selection do not satisfy the red slope criterion, with the majority originating from CONGRESS. In these cases, the [O III] doublet falls within the F277W filter, making the F277W F444W color less red, and causing these sources to fall outside the selection boundary. An example of an object that clearly exhibits LRD-like behavior in its SED but is not color-selected as an LRD due to strong [O III] emission is presented in Rinaldi et al. (2025). This further highlights the challenges and limitations of color-based LRD selection at these redshifts.
@@
While several alternative color selection criteria of LRDs have been proposed in the literature (e.g., Greene et al. 2024; Pérez-González et al. 2024), we adopt the criteria used by Rinaldi et al. (2024) in this work to maintain consistency, as their LRD sample is also constructed in the GOODS-N field.
@@
To further explore the relationship between broad-line emitters and photometrically selected LRDs, we perform a comparison between the average SEDs of our broad H emitters and the LRD sample selected in Rinaldi et al. (2024). We follow the methodology outlined in Pérez-González et al. (2024) to derive the average SED of the LRD sample by stacking the SEDs of LRDs with available spectroscopic redshifts (including their emission lines) and computing the mean flux for every ten consecutive data points in the wavelength-ordered list. In contrast, the average SED of our broad H emitter sample is constructed using the best-fit SEDs obtained in Section V.3, as about half of the sources in our sample have photometry available in only 6 filters. Figure 10 displays the average rest-frame SEDs for both broad H emitters and LRDs, with flux normalized at 0.4 m. The average SED of LRDs selected in the GOODS-S field presented in Pérez-González et al. (2024) is also included for comparison.
@@
Overall, we do not observe significant differences between the average SEDs of our broad H emitters and LRDs within the rest-frame 0.1–1 m range. The SEDs closely align up to rest-frame 0.4 m, suggesting that LRDs may share similar UV emission properties with broad-line emitters. However, beyond 0.5 m, the average SEDs of LRDs exhibit a noticeably steeper slope toward longer wavelengths compared to those of broad-line emitters, with particularly significant deviations at 0.5 and 0.65 m. Given that the observed continuum in the average LRD SEDs between these wavelengths is comparable to that of broad-line emitters, we suspect that the steep red slope is not an intrinsic feature of at least some LRDs, but rather a result of line-boosting effects from strong emission lines, such as [O III] and H, falling within individual wide filters (Hainline et al., 2025). Figure 10 further supports the concern that adopting a low red slope cut for LRD selection may lead to contamination by sources that exhibit “fake” steep red slopes induced by such line-boosting effects, particularly when there are strong emission lines falling within the filters used for color selection.
V.5 Properties of Non-LRD AGN Galaxies
It should be noted that BLAGN galaxies that do not satisfy the NIRCam-based red-slope criterion (and therefore drop out the LRD selection) could still exhibit an SED upturn (i.e., a V-shaped SED) at 5 µm (i.e., beyond the NIRCam wavelength coverage). A good example is JADES 1089568 at 4.05 shown in the bottom panel of Figure 7. The figure shows that this object has a very blue color over the NIRCam passbands (e.g., due to its star-forming host), and in fact it is one of the bluest objects in Figure 9 in terms of the F277W-F444W color and slope. And yet it clearly exhibits a red SED rising beyond 5 µm ( 1 µm), which can be fit well with our AGN SED model. Such an SED is actually typical of AGN galaxies and widely seen across the redshift (e.g., Lyu et al., 2024) with hot dust emission from the AGN torus starting to dominate only in the rest-frame near-/mid-infrared. This illustrates why the LRD selection (and more generally NIRCam-only color selections) is highly incomplete when selecting AGN galaxies (see also Hainline et al. (2025)).
V.6 Detectability of Broad H Line in Grism Spectra
@@
We identify 19 photometrically selected LRDs from Rinaldi et al. (2024) that lie within the FRESCO and CONGRESS fields and have H emission detected in the grism spectra. However, 68% of these sources (13 out of 19) do not show broad H line profiles (i.e.,FWHM 1000 km/s) in the grism spectroscopy, with the exception of the 6 broad H emitters identified in this work. Such a low fraction of LRDs displaying broad-line features is consistent with the findings reported by Pérez-González et al. (2024).
@@
Although no broad-line feature is observed in these LRDs, we cannot fully rule out the possibility that they are broad-line emitters, as the broad component could be hidden due to the insufficient sensitivity of the grism spectra. Notably, two sources that show no broad line feature in the grism spectra have been reported to exhibit broad line features in NIRSpec medium-resolution grating data in previous work (Maiolino et al., 2024). To illustrate how S/N may affect the detection of broad components, we present both the NIRSpec/grating and NIRCam/grism spectra for these two sources in Appendix C.
@@
Figure 11 presents the distribution of our broad-line emitters and the photometrically selected LRDs of Rinaldi et al. (2024), which lack broad-line features in the grism spectra, on the redshift-magnitude plane. Overall, our broad-line emitters are significantly brighter than the LRDs, with a division at 26 mag that effectively separates the two populations. Notably, the two sources that exhibit broad-line features in the grating spectra but not in the grism are both fainter than 26 mag, suggesting that some faint LRDs may indeed be broad-line emitters, but their broad components are too faint to be detected in grism observations. In other words, the broad-line search with the CONGRESS/FRESCO data is effective only down to 26 mag.
@@ VI Conclusion
Among the 19 broad H emitters, 18 of them have a broad component that contributes more than 50% of the total H flux. Additionally, [O III] emission line is detected in three sources, with line widths significantly narrower than those of the broad H components, suggesting that the observed broad features of these sources are mainly driven by the AGN BLR. By computing the broad H LF of our sample, we find that it is consistent with those of other JWST-selected BLAGN in the literature.
@@
The stellar masses of the host galaxies in our sample were derived through SED-fitting using Prospector. Consistent with previous studies, our sample exhibits higher ratios relative to the local relations reported by Reines and Volonteri (2015) and Greene et al. (2020). However, given that our sample is selected based on the detection of broad H emission, some degree of selection bias may be present. In particular, low sources may be missed due to the limited sensitivity of the grism, which could prevent the detection of faint broad-line components.
By testing various LRD selection techniques proposed in the literature, we find that all of these methods will miss a significant fraction of our spectroscopically confirmed broad H emitters, while the majority of the broad H emitters that failed the LRD color selection lack the steep red slopes. This suggests that not all broad-line emitters with compact and red appearance exhibit the characteristic “V-shaped” SEDs of LRDs.
To further investigate the nature of LRDs, we compared the average SED of our broad-line emitter sample with that of the photometrically selected LRDs from Rinaldi et al. (2024) in the same field. The average SEDs are largely consistent with each other, except for a steeper red slope in the LRD sample at rest-frame wavelength 0.5 m, likely due to the contribution of [O III] and H emission. Given the much lower continuum levels observed between these two lines, we conclude that the red slope, at least in some LRDs, would be naturally less steep if emission lines are removed, as discussed in Hainline et al. (2025).
We further examined the LRDs in the GOODS-N field with H detections in the grism spectra and find that 68% do not exhibit broad H features. This suggests that the LRD population may not only consist of broad-line emitters. Additionally, we identified two LRDs that lack broad H features in the NIRCam grism spectra but do show broad components in the NIRSpec grating spectra. Both of these sources are fainter than 26 mag in the filter where H is observed, whereas the majority of our broad H emitters are brighter than this threshold. This implies that the broad-line search with the CONGRESS/FRESCO data is effective only down to 26 mag, and therefore that some LRDs may indeed host broad-line components that are undetected in the grism data due to its limited sensitivity.
@@ Acknowledgments
EE, YZ acknowledge support from the NIRCam Science Team contract to the University of Arizona, NAS5-02015. AJB acknowledges funding from the “FirstGalaxies” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 789056). RM acknowledges support by the Science and Technology Facilities Council (STFC), by the ERC through Advanced Grant 695671 “QUENCH”, and by the UKRI Frontier Research grant RISEandFALL. RM also acknowledges funding from a research professorship from the Royal Society. BER acknowledges support from the NIRCam Science Team contract to the University of Arizona, NAS5-02015, and JWST Program 3215. ST acknowledges support by the Royal Society Research Grant G125142. GV acknowledges support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. The research of CCW is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
@@
This work is based on observations made with the NASA/ESA Hubble Space Telescope and NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1181 (JADES), 1895 (FRESCO), and 3577 (CONGRESS). The specific JWST observations analyzed can be accessed via http://dx.doi.org/10.17909/6rfk-6s81 (catalog doi:10.17909/6rfk-6s81 ). Support for program #3577 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. The authors acknowledge the FRESCO team for developing their observing program with a zero-exclusive-access period.
Appendix A Non-Confirmation of Two Broad H Emitters Reported by Covelo-Paz et al. (2024)
@@
In our sample, we exclude two broad H emitters reported by Covelo-Paz et al. (2024) due to concerns regarding the validity of their classification. For the first source (JADES ID: 1080661; RA = 189.27616 deg, Dec = 62.21416 deg), the flux ratio does not meet our compactness criterion. Given that the selection criteria in Covelo-Paz et al. (2024) do not incorporate compactness, this mismatch is natural. Figure 12 presents the 1D and 2D grism spectra obtained in F356W from CONGRESS. The F356W direct image reveals an extended morphology for this source. We also performed a 1D line fit of its H line profile. As illustrated in Figure 13, even after accounting for potential morphological broadening, the best-fit FWHM is only 500 km/s, which does not satisfy our requirement for the line width of a broad H emitter. Additionally, we note that our best-fit FWHM differs from the value reported in Covelo-Paz et al. (2024) (i.e., 1674 km/s). However, the cause of this discrepancy is unclear.
@@
For the second source (JADES ID: 1089616; RA = 189.18713 deg, Dec = 62.27289 deg), the top panel of Figure 14 shows its 1D and 2D grism spectra in F356W from CONGRESS. While a strong emission feature is detected at 3.8 m, we observe a noticeable spatial offset between the position of the emission line and that of the source in the direct image, raising concerns about its association with the target. In the bottom panel of Figure 14, which presents the F444W grism spectra from FRESCO for the same source, an emission line at 4.15 m is detected with no spatial offset and a S/N4. Combined with the detection of another emission line at 3.17 m, we identify these features as H and [O III] of this source, respectively, confirming the redshift of the source is 5.324. The relatively weak line strengths are also consistent with the faint nature of the source in the direct images. Upon further inspection, the strong 3.8 m line is confirmed to be [S III] from a nearby bright source (JADES ID: 1029833; RA = 189.20103 deg, Dec = 62.26567 deg). Figure 15 presents the 1D and 2D grism spectra of this source in the F356W and F444W filters. The [S III] emission line of this source matches the 3.8 m emission line observed in the 2D spectrum of source 1089616 very well, suggesting the 3.8 m emission line is not H of 1089616, but [S III] of 1029833.
@@
Appendix B Discrepant Broad-Line Widths for a BLAGN Reported by Greene et al. (2024)
@@
Recently, the JWST Cycle 3 General Observer program Slitless Areal Pure-Parallel High-Redshift Emission Survey (SAPPHIRES-PID: 6434; PI: E. Egami; Sun et al. 2025a) revisited the Abell 2744 field and obtained NIRCam grism spectrum (R1600) for one of the BLAGN (Prism ID: 38108; SAPPHIRES’s ID: 39082) identified in Greene et al. (2024). Figure 16 shows the observed H line in 1D and 2D grism spectra obtained in F444W filter, along with the line fitting results. The 1D grism spectrum reveals a broad H component with a line width of 1300 km/s, which contrasts significantly with the prism-based value of 4000 km/s. For comparison, we also include the H broad component observed in prism by using the FWHM and H flux (treated as an upper limit of the flux for the broad component) reported by Greene et al. (2024). Even a modest difference in the FWHM measurements can result in a substantial variation in the inferred . In Figure 6, we plot the AGN bolometric luminosity and black hole mass derived for this source from the SAPPHIRES NIRCam grism data (the “X” marker). With the smaller inferred from the grism spectrum, the source becomes more consistent with the distribution of our sample, providing a potential explanation for the discrepancy in black hole mass estimates.
@@
However, the underlying cause of the discrepancy in FWHM measurements remains unclear. Taking into account the large uncertainty associated with the prism-based FWHM, our measurement and that from Greene et al. (2024) are broadly consistent with each other at the 2- level. Since our line fitting was performed on spectra without continuum subtraction, we can rule out the possibility that the broad component was lost due to over-subtraction of continuum. There are two possible explanations for this discrepancy. First, the difference in FWHM values may arise from differences in the respective LSFs. Second, the grism observation may lack the sensitivity required to detect such a faint broad component. Furthermore, we note that Greene et al. (2024) identified broad-line sources using slightly different selection criteria, in particular requiring the FWHM of the broad H component to 2000 km s-1. This difference in criteria may also contribute to the discrepancy in FWHM measurements.
@@
In this work, we identify two photometrically selected LRDs from Rinaldi et al. (2024) that exhibit broad H line profiles in the NIRSpec medium resolution grating spectra presented by Maiolino et al. (2024), but not in the NIRCam grism spectra, likely due to the limited sensitivity of the grism observations. Figure 17 illustrates the differences in the observed H line profiles for these two sources between the grism and grating data. The S/N of the grating spectra are slightly lower than those presented in Maiolino et al. (2024), possibly due to differences in the spectra extraction methods. In both cases, the H lines appear significantly narrower in the grism spectra compared to the grating spectra. Additionally, the grism continuum exhibits a much higher noise level, supporting the possibility that the broad components are undetected due to the lower S/N of the grism observations.